Circular No. 7962

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 Q5 (LINEAR) L. Manguso, Lincoln Laboratory, Massachusetts Institute of Technol-

ogy, reports the discovery of a comet (discovery observation given below) on LINEAR images. Following posting on the NEO Confirmation Page, the object’s cometary nature was confirmed by several CCD observers. J. Nomen (0.40-m Schmidt telescope, Ametlla de Mar, Spain) reported a 20′′ coma (m_{1} = 16.7) and a 1′ tail in p.a. 140^{o} on Aug. 29.0 UT. Near the same time, P. Kušnirak (0.65-m f/3.6 reflector, Ondřejov) found the comet to be slightly diffuse with a coma diameter of 10′′ and a faint tail toward the southeast on R-band images. On Aug. 29.3, P. R. Holvorcem and M. Schwartz (0.36-m reflector, Cottage Grove, OR) found a coma diameter of ∼ 20′′ and m_{1} = 15.6 on 90-s exposures.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 28.21840  20^{h}06^{m}07.^{s}55  +38^{o}24′57′′.5  17.3

The available astrometry, the following preliminary parabolic orbital elements (from 25 observations, Aug. 28–29), and an ephemeris appear on MPEC 2002-Q43.

T = 2002 Nov. 19.957 TT  ω = 135.^{o}275
Ω = 34.322 }2000.0
q = 1.22115 AU  i = 149.318
2000 CF_{105}

W. Romanishin, University of Oklahoma; S. Tegler, Northern Arizona University; K. Noll and D. Stephens, Space Telescope Science Institute; W. Grundy, J. Spencer, R. Millis, and M. Buie, Lowell Observatory; and D. Cruikshank, Ames Research Center, confirm the binary nature of the transneptunian object 2000 CF_{105}. The image of the object appears elongated in the coaddition of four 10-min R-band exposures taken (under 0′′.7 seeing conditions) with the Keck I 10-m telescope (+ LRIS) on Apr. 11.3 UT. Preliminary image modeling shows a separation of 0′′.8 ± 0′′.2 at p.a. 103^{o} ± 5^{o}, with a difference of 0.6 ± 0.2 magnitude between the components. These numbers are consistent with (but have larger uncertainties than) the numbers reported for the Hubble Space Telescope observations on Jan. 12.1 (cf. IAUC 7857).

2002 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green